OpenStudy (anonymous):

If sin y+cos y=x verify the identity (cos y-sin y)^3=(2-x^2)^3/2

OpenStudy (anonymous):

given sin y+cos y=x square both sides and find the value of 2siny cos y

OpenStudy (ybarrap):

Start with \(~\sin y+\cos y=x\) Square both sides: \((\sin y+\cos y)^2=\sin^2y+2\cos y\sin y+\cos^2=x^2\\\) Next, given \((\cos y-\sin y)^3=(2-x^2)^{3/2}\\\), take cube roots on both sides: \(\cos y-\sin y=(2-x^2)^{1/2}\\\) Now square both sides:\((\cos y-\sin y)^{2}=2-x^2\\\) Substitute \(x^2\) derived above and simplify: \(\cos^2y-2\sin y\cos y+\cos^2y=2-\sin^2y-2\cos y\sin y-\cos^2\) \( 1-2\sin y\cos y=2-2\cos y\sin y-1\\ -2\sin y\cos y=-2\cos y\sin y \) Which shows that \(\sin y+\cos y=x\) is a solution to \((\cos y-\sin y)^3=(2-x^2)^{3/2}\) .